OpenStudy (anonymous):

An equilateral triangle has a perimeter of 48 cm. What is its area?

OpenStudy (anonymous):

each side is 16 cm, which is the length of your base. Draw out the diagram, and use Pythagorean theorem to see your height= sqrt(16^2-8^2)= 8sqrt(3) A=1/2bh=16*8sqrt(3)/3

OpenStudy (anonymous):

so it is 16 times 8

OpenStudy (anonymous):

no your final answer should be 64*sqrt(3)=111

OpenStudy (anonymous):

so it would eqaul 111

OpenStudy (anonymous):

yes your area is 111 cm2

OpenStudy (anonymous):

do i use square root for everything im finfing the area for

OpenStudy (anonymous):

@jdilorenzo3

OpenStudy (anonymous):

 the answer would be 18 right